# Fast3R: Towards 3D Reconstruction of 1000+ Images in One Forward Pass

### v1

@@ 1 Introduction
We propose Fast3R, a novel multi-view reconstruction framework designed to overcome these limitations. Building on DUSt3R’s foundations, Fast3R leverages a Transformer-based architecture [56] that processes multiple images in parallel, allowing $N$ images to be reconstructed in a single forward pass. By eliminating the need for sequential or pairwise processing, each frame can simultaneously attend to all other frames in the input set during reconstruction, significantly reduces error accumulation. Perhaps surprisingly, Fast3R also takes significantly less time.
@@ 2 Related Work
Multi-view 3D reconstruction: Almost all modern 3D reconstruction approaches are based on the traditional multi-view geometry (MVG) pipeline [21]. MVG-based methods first identify corresponding pixels between image pairs, and then use camera models and projective multiview geometry to lift these correspondences to 3D points. The process happens in sequential stages: feature extraction, finding pairwise image correspondences, triangulation to 3D and pairwise relative camera pose, and global bundle alignment. However, any pipeline approach is prone to accumulating errors, which are especially common in the hand-crafted components. Moreover, the sequential nature prevents parallelization, which limits speed and scalability.
@@
Pointmap representation: DUSt3R [61] takes this evolution the furthest by proposing pointmap regression to replace everything in the MVG pipeline up to global pairwise alignment. Rather than first attempting to solve for camera parameters in order to triangulate corresponding pixels, DUSt3R trains a model to directly predict 3D pointmaps for pairs of images in a shared coordinate system. Other MVG component tasks such as relative camera pose estimation and depth estimation can be recovered from the resulting pointmap representation. However, DUSt3R’s pairwise assumption is a limitation, as it requires inference on $\mathcal{O}(N^{2})$ image pairs and then a global alignment optimization, which is per-scene and does not improve with more data. Moreover, this process quickly becomes slow or crashes due to exceeded system memory, even for relatively modest numbers of images.
@@ 3.1 Problem definition
Fast3R predicts local and global pointmaps $\mathbf{X}_{L}$ and $\mathbf{X}_{G}$, and corresponding confidence maps $\Sigma_{L}$ and $\Sigma_{G}$ (of shape $\Sigma\in\mathbb{R}^{N\times H\times W}$). Fast3R maps $N$ RGB images to
@@
Like MASt3R, the global pointmap $\mathbf{X}_{G}$ is in the coordinate frame of the first camera and the $\mathbf{X}_{L}$ is in the coordinate frame of the viewing camera, as shown in Figure 2
@@ 3.3 Model architecture
Fusion transformer: Most of the computation in Fast3R happens in the fusion transformer, which has a generic architecture. We use a 12-layer transformer similar to ViT-B [12] or BERT [10], however this could be scaled up. This fusion transformer takes the concatenated encoded image patches from all views and performs all-to-all self-attention. This operation provides Fast3R with full context, beyond the information provided in pairs alone.
Pointmap head: Finally, Fast3R uses separate DPT-L [37] decoder heads to map these tokens to the local and global pointmaps $(\mathbf{X}_{\text{L}},\mathbf{X}_{\text{G}})$, and confidence maps $(\Sigma_{L},\Sigma_{G}$).
Image index positional embedding generalization: We would like Fast3R to be able to handle many views at inference, more than were used to train a model. A naïve way to embed views during testing would be to embed them in the same way as training: i.e. use the same Spherical Harmonic frequencies  [49] to embed raw indices $SH(\{1,...,N\})$ during training, and $SH(\{1,...,N_{\mathrm{test}}\})$ during inference. In LLMs this causes poor performance, and in preliminary experiments we also found that the resulting model did not work well when the number of input images exceeded that used during training. We therefore adapt Position Interpolation [5], a solution from LLMs, where during training we randomly draw $N$ indexes from a larger pool $N^{\prime}$ of possible samples. [5] draws samples using a regular grid since the LLM inputs form a regular ordered sequence. Our images are unordered, so we draw $N\subset\{1,...,N^{\prime}\}$ uniformly at random. To the transformer, the strategy looks indistinguishable from masking out images, and $N^{\prime}\gg N$ controls the masking ratio. [^2] Patches $H_{1}$ from the first image $I_{1}$ are always embedded with $p_{1}$, since that image provides the coordinate frame for the global head. This strategy enables Fast3r to handle $N=1000$ images during inference, even if only trained with $N=20$ images.
@@ 3.4 Memory-Efficient Implementation
For example, model size and throughput can be increased by sharding the model and/or data minibatch across multiple machines, such as through model parallelism [45, 22], data parallelism [26], and tensor parallelism [46, 32]. During training, optimizer weights, states, and gradients can also be sharded [35]. Systems-level advances have also been proposed, such as FlashAttention [8, 7], which uses GPU kernels leveraging the hardware topology to compute attention in a time and memory-efficient way. These are implemented in libraries such as FAIRScale [16], DeepSpeed [35] and Huggingface [64], and require significant engineering effort.
@@
Methods Co3Dv2 [39] FPS
RRA@15$\uparrow$ RRA@5$\uparrow$ RTA@15$\uparrow$ RTA@5$\uparrow$ mAA(30$\uparrow$
Colmap+SG [9, 41] 36.1 24.4 27.3 17.2 25.3 0.056
PixSfM [28] 33.7 26.1 32.9 17.6 30.1 -
RelPose [70] 57.1 - - - - 0.02
PosReg [59] 53.2 - 49.1 - 45.0 0.015
PoseDiff [59] 80.5 59.5 79.8 61.7 66.5 0.015
RelPose++ [27] (85.5) - - - - 0.02
DUSt3R [60] 96.2 - 86.8 - 76.7 0.78
MASt3R [24] 94.6 93.2 91.9 86.2 81.8 0.23
Fast3R (Ours) 99.7 97.4 87.1 76.1 82.5 251.1
Table 1: Multi-view pose regression on the CO3Dv2 [39] dataset with 10 random frames. Parentheses denote methods that do not report results on the 10 views set; we report their best for comparison (8 views). Fast3R does not assume known camera intrinsics.
@@ 4 Experiments
Training Data We train on a mix of real-world object-centric and scene scan data: CO3D [39], ScanNet++ [67], ARKitScenes [4], and Habitat [43]. We use a subset of the datasets in DUSt3R, specifically 4 of the 9 datasets, for a total of around 2000 unique scenes scans and 1300 videos of 50 object classes. Note that this is only 7% of CO3D, which is also what the baselines DUSt3R [61], Spann3R [57], and MASt3R [25] use.
Baselines DUSt3R [61] is the closest approach to ours, and competitive on visual odometry and reconstruction benchmarks. That paper contains extensive comparisons against other methods, and we adopt it as our main baseline. We additionally consider DUSt3R’s follow-up work, MASt3R [25], as well as a concurrent work Spann3R [57], which also seeks to replace DUSt3R’s expensive global alignment stage by sequentially processing frames with a spatial memory. For camera pose estimation and 3D reconstruction, we include comparisons to task-specific methods.
Architecture Details In our experiments we use the following components for the meta-architecture:
@@
    The Image Encoder uses a CroCo ViT-B [63] architecture, initialized with DUSt3R pretrained weights [61].
@@
    The Fusion Transformer is a 12-layer ViT-B model using the BERT architecture with 12 layers, 12 heads, embedding dimension size 768 and MLP ratio 4.0. We use a pool size of $N^{\prime}=1000$ image index embeddings.
@@
    The Pointmap Decoder is a dense vision Transformer following the DPT-L architecture  [37].
Training Details Our models are trained on $512\times 512$ images using AdamW [29] for 6.5K steps, with a learning rate of 0.0001 and cosine annealing schedule. Unlike DUSt3R, we do not use staged training. For the most part, we use the same dataloaders used in the public baseline implementations. We train with batch size 64, with each sample consisting of a tuple of $N=20$ views. This process takes 6.13 days on 64 A100 GPUs. We additionally make use of several strategies to enable efficient training. First, we use the FlashAttention [8, 7] to improve time and memory efficiency. Even so, a naïve implementation runs out of memory even with batch size 1 when $N>16$, so we use DeepSpeed ZeRO stage 2 training [35], whereby optimizer states, moment estimates, and gradients are partitioned across different machines. This enables us to train with up to $N=28$ views per data sample, with a batch size of one per GPU.
@@ 4.1 Inference Efficiency
Table 3: Quantitative reconstruction results on scene datasets: The numbers indicate median distance to GT points on 7-Scenes [47] and NRGBD [3] datasets. These datasets contain video trajectories of 500-1500 frames, and we evaluate using the same frame skip as other baselines. For 7-Scenes this is skip=20, and NRGBD uses skip=40. DUSt3R indicates using DUSt3R’s final weights on $224\times 224$ images, to fit within GPU memory. Distances are scaled 100×100$\times$100 × to remove the leading 0.0. We defer the full table to the appendix.
@@ 4.2 Camera Pose Estimation
We report Relative Rotation Accuracy (RRA) and Relative Translation Accuracy (RTA) at a threshold of $15^{\circ}$, mean Average Accuracy (mAA) at threshold $30^{\circ}$, and model frames per second (FPS) in Table 1. On Co3D, Fast3R surpasses all other methods across the RRA and mAA metrics, achieving near-perfect RRA, while remaining competitive on RTA. Importantly, it is also orders of magnitude faster: $\mathbf{200\times}$ faster than DUSt3R and $\mathbf{700\times}$ faster than MASt3R.
@@
We found that local pointmaps learn finer detail than the global pointmaps. Therefore we use the local pointmaps for detail and the global pointmaps for the high-level structure. Specifically, we independently align each image’s local pointmap to the global pointmap using ICP, and use aligned local pointmaps for evaluation.
@@ 5.1 Scaling the number of views
Fast3R is able to use all-to-all attention during training, which lets it learn from global context. We hypothesize that the additional context provided by more views during training allows the model to learn higher-order correspondences between multiple frames, ultimately increasing model performance and increasing potential for scaling.
Figures 7 and 8 shows that training on increasingly more views consistently improves RRA and RTA for visual odometry, and reconstruction accuracy—even when the number of views used during evaluation is held constant and the model is ultimately evaluated on fewer views than were seen during training. We further evaluate the model’s ability to reason about additional views by increasing the number of images that Fast3R sees during inference. Figure 4 and Figure 5 indicate that as the model uses more views, the average per-view performance improves. This behavior holds for all evaluated metrics in both camera pose estimation and reconstruction. As shown in Figure 5, the model has a better per-view accuracy using 50 images than it does with 20, even though it was trained with 20. Many applications (e.g. reconstruction, odometry) require inference on many views, which is a major motivation for Fast3R removing the pairwise constraint.
@@ 5.2 Training without position interpolation
In Section 3.3, we introduced a randomized version of [5] to enable inference on more views than seen training. Without this technique, model accuracy quickly degrades for pointmap corresponding to image indexes outside the training range, as shown in Figure 9 (top). A version of Fast3R trained on $N=4$ views still produces high quality pointmaps for views in slot 5 to 24 (Figure 9 bottom).
5.3 Removing the local decoder head
Table 5 shows that removing the local head learns finer details before the global head does. We hypothesize that this is because the global head needs to first learn the coordinate system and then learn the fine details. As the model improves, the local head may not be necessary.
Method 7-Scenes [47] NRGBD [3] DTU
@@ 5.3 Removing the local decoder head
Fast3R (ours) 2.84 1.37 4.39 1.28 3.91 1.41
w/o local head 4.81 1.64 4.85 1.32 3.88 1.41
@@ 6 Conclusion
The architecture of Fast3R allows for parallel processing of many views, and its positional embedding design enables “train short, test long” in terms of context length of views. However, we observed that for scenes with very large reconstruction areas, when the number of views becomes extreme (e.g., more than 200), the point map of some views (particularly those with a low confidence score) begins to exhibit drifting behavior. One current way to address this issue is to drop frames with low confidence scores. In dense reconstruction, this approach typically does not hurt reconstruction quality too much. However, to fundamentally address this problem, we hypothesize that future work could explore the following avenues: (1) incorporating more data containing large scenes to improve generalization to such cases; (2) designing better positional embeddings inspired by state-of-the-art long-context language models [74], which can handle very long context lengths and exploit the temporal structure of ordered image sequences (e.g., video).
@@
Figure 10: Visualizations of results on NRGBD scenes. Fast3R learns the regularity of indoor rooms (square-like shapes) and demonstrates “loop closure” capabilities.

### v2

@@ 1 Introduction
We propose Fast3R, a novel multi-view reconstruction framework designed to overcome these limitations. Building on DUSt3R’s foundations, Fast3R leverages a Transformer-based architecture [56] that processes multiple images in parallel, allowing $N$ images to be reconstructed in a single forward pass. By eliminating the need for sequential or pairwise processing, each frame can simultaneously attend to all other frames in the input set during reconstruction, significantly reducing error accumulation. Perhaps surprisingly, Fast3R also takes significantly less time.
@@ 2 Related Work
Multi-view 3D reconstruction: Almost all modern 3D reconstruction approaches are based on the traditional multi-view geometry (MVG) pipeline [20]. MVG-based methods first identify corresponding pixels between image pairs, and then use camera models and projective multiview geometry to lift these correspondences to 3D points. The process happens in sequential stages: feature extraction, finding pairwise image correspondences, triangulation to 3D and pairwise relative camera pose, and global bundle alignment. However, any pipeline approach is prone to accumulating errors, which are especially common in hand-crafted components. Moreover, the sequential nature prevents parallelization, which limits speed and scalability.
@@
Pointmap representation: DUSt3R [61] takes this evolution the furthest by proposing pointmap regression to replace everything in the MVG pipeline up to global pairwise alignment. Rather than first attempting to solve for camera parameters in order to triangulate corresponding pixels, DUSt3R trains a model to directly predict 3D pointmaps for pairs of images in a shared coordinate frame. Other MVG component tasks such as relative camera pose estimation and depth estimation can be recovered from the resulting pointmap representation. However, DUSt3R’s pairwise assumption is a limitation, as it requires inference on $\mathcal{O}(N^{2})$ image pairs and then a global alignment optimization, which is per-scene and does not improve with more data. Moreover, this process quickly becomes slow or crashes due to exceeded system memory, even for relatively modest numbers of images.
@@ 3.1 Problem definition
Fast3R predicts two pointmaps: local pointmap $\mathbf{X}_{L}$ and global pointmap $\mathbf{X}_{G}$, and corresponding confidence maps $\Sigma_{L}$ and $\Sigma_{G}$ (of shape $\Sigma\in\mathbb{R}^{N\times H\times W}$). Overall, Fast3R maps $N$ RGB images to $N$ local and global pointmaps:
@@
The global pointmap $\mathbf{X}_{G}$ is in the coordinate frame of the first camera and the $\mathbf{X}_{L}$ is in the coordinate frame of the viewing camera, as shown in Figure 2
@@ 3.3 Model architecture
Fusion Transformer: Most of the computation in Fast3R happens in the fusion transformer. We use a 24-layer transformer similar to ViT-L [11]. This fusion transformer takes the concatenated encoded image patches from all views and performs all-to-all self-attention. This operation provides Fast3R with full context from all views, beyond the information provided in pairs alone.
Pointmap Decoding Heads: Finally, Fast3R uses two separate DPT [37] decoding heads to map these tokens to local and global pointmaps $(\mathbf{X}_{\text{L}},\mathbf{X}_{\text{G}})$, and confidence maps $(\Sigma_{L},\Sigma_{G}$).
Image index positional embedding generalization: We would like Fast3R to be able to handle many views at inference, more than were used to train a model. A naïve way to embed views during testing would be to embed them in the same way as training: i.e. use the same Spherical Harmonic frequencies  [49] to embed raw indices $SH(\{1,...,N\})$ during training, and $SH(\{1,...,N_{\mathrm{test}}\})$ during inference. In LLMs this causes poor performance, and in preliminary experiments, we also found that the resulting model did not work well when the number of input images exceeded that used during training (Sec. 5.3). We therefore adopt Position Interpolation [5], a solution from LLMs, where during training we randomly draw $N$ indexes from a larger pool $N^{\prime}$ of possible samples. [5] draws samples using a regular grid since the LLM inputs form a regular ordered sequence. Our images are unordered, so we draw $N\subset\{1,...,N^{\prime}\}$ uniformly at random. To the transformer, the strategy looks indistinguishable from masking out images, and $N^{\prime}\gg N$ controls the masking ratio. [^2] Patches $H_{1}$ from the first image $I_{1}$ are always embedded with $p_{1}$, since that image provides the coordinate frame for the global head. This strategy enables Fast3R to handle $N=1000$ images during inference, even if only trained with $N=20$ images.
@@ 3.4 Memory-Efficient Implementation
For example, model size and throughput can be increased by sharding the model and/or data minibatch across multiple machines, such as through model parallelism [45, 21], data parallelism [25], and tensor parallelism [46, 32]. During training, optimizer weights, states, and gradients can also be sharded [35]. Systems-level advances have also been proposed, such as FlashAttention [8, 7], which uses highly optimized GPU kernels leveraging the hardware topology to compute attention in a time and memory-efficient way. These are implemented in libraries such as FAIRScale [15], DeepSpeed [35] and HuggingFace [64], and require significant engineering effort.
@@
Methods CO3Dv2 [39] RealEstate10K [75] FPS
RRA@15$\uparrow$ RRA@5$\uparrow$ RTA@15$\uparrow$ RTA@5$\uparrow$ mAA(30$\uparrow$ mAA(30$\uparrow$
Colmap+SG [9, 41] 36.1 24.4 27.3 17.2 25.3 45.2 0.056
PixSfM [28] 33.7 26.1 32.9 17.6 30.1 49.4 -
RelPose [71] 57.1 - - - - - 0.02
PosReg [59] 53.2 - 49.1 - 45.0 - 0.015
PoseDiff [59] 80.5 59.5 79.8 61.7 66.5 48.0 0.015
RelPose++ [27] (85.5) - - - - - 0.02
DUSt3R [60] 96.2 - 86.8 - 76.7 67.7 0.78
MASt3R [23] 94.6 93.2 91.9 86.2 81.8 76.4 0.23
Fast3R-no-outdoor (Ours) 99.7 97.4 87.1 76.1 82.5 - 251.1
Fast3R (Ours) 96.2 90.2 81.6 68.2 75.0 72.7 251.1
Table 1: Multi-view pose regression on the CO3D [39] and RealEstate10K [75] datasets. Parentheses denote methods that do not report results on the 10 views set; we report their best for comparison (8 views). Fast3R does not assume known camera intrinsics.
@@ 4 Experiments
Training Data We train on a mix of real-world object-centric and scene scan data: CO3D [39], ScanNet++ [68], ARKitScenes [4], Habitat [43], BlendedMVS [67], and MegaDepth [26]. This is a subset of the datasets in DUSt3R, specifically 6 of the 9 datasets.
Baselines DUSt3R [61] is the closest approach to ours, and competitive on visual odometry and reconstruction benchmarks. That paper contains extensive comparisons against other methods, and we adopt it as our main baseline. We additionally consider DUSt3R’s follow-up work, MASt3R [24], as well as a concurrent work Spann3R [57], which also seeks to replace DUSt3R’s expensive global alignment stage by sequentially processing frames with an external spatial memory. For camera pose estimation and 3D reconstruction, we include comparisons to task-specific methods.
Architecture Details In our experiments, we use the following components for the meta-architecture:
@@
    The Image Encoder uses a ViT-Large [63] architecture, initialized with DUSt3R pretrained weights [61]. The ViT-L uses 16x16 patch size, and has 24 layers, 16 heads, embedding dimension size 1024, and MLP ratio 4.0.
@@
    The Fusion Transformer is a ViT-Large model initialized from scratch. We use a pool size of $N^{\prime}=1000$ for image index embeddings.
@@
    The Pointmap Decoding Heads include two heads: global and local, both of which are DPT heads following [37]. We initialize the global head with DUSt3R pretrained weights, and initialize the local head from scratch.
Training Details Our models are trained on images with 512 resolution (512 on the longest side) using AdamW [29] for 174K steps, with a learning rate of 0.0001 and cosine annealing schedule. Unlike DUSt3R, we do not use staged training. We implement multi-view dataloaders that can load N images in each sample. We train with batch size 128, with each sample consisting of a tuple of $N=20$ views. This process takes 6.13 days on 128 Nvidia A100-80GB GPUs. We additionally make use of several strategies to enable efficient training. First, we use the FlashAttention [8, 7] to improve time and memory efficiency. Even so, a naïve implementation runs out of memory even with batch size 1 when $N>16$, so we use DeepSpeed ZeRO stage 2 training [35], whereby optimizer states, moment estimates, and gradients are partitioned across different machines. This enables us to train with up to $N=28$ views per data sample at max, with a batch size of one per GPU.
@@ 4.1 Inference Efficiency
Table 3: Quantitative reconstruction results on scene datasets: The numbers indicate median distance to GT points on 7-Scenes [47] and NRGBD [3] datasets. These datasets contain video trajectories of 500-1500 frames, and we evaluate using the same frame skip as other baselines. For 7-Scenes this is skip=20, and NRGBD uses skip=40. DUSt3R indicates using DUSt3R’s final weights on $224\times 224$ images, to fit within GPU memory. Distances are scaled 100×100$\times$100 × to remove the leading 0.00.
@@ 4.2 Camera Pose Estimation
We report Relative Rotation Accuracy (RRA) and Relative Translation Accuracy (RTA) at a threshold of $15^{\circ}$, mean Average Accuracy (mAA) at threshold $30^{\circ}$, and model frames per second (FPS) in Table 1. On CO3D, Fast3R surpasses all other methods across the RRA and mAA metrics, achieving near-perfect RRA, while remaining competitive on RTA. Importantly, it is also orders of magnitude faster: $\mathbf{320\times}$ faster than DUSt3R and $\mathbf{1000\times}$ faster than MASt3R. Fast3R-no-outdoor is the same Fast3R model but trained without the BlendedMVS and MegaDepth datasets. We find these datasets help the model generalize to more diverse scenes (e.g., drone shots and large outdoor landmarks) but slightly hurt pose estimation performance on CO3D.
@@
We found that local pointmap head learns more accurate pointmaps than the global pointmap head (ablation in Sec. 5.4). Therefore we use the local pointmaps for detail and the global pointmaps for the high-level structure. Specifically, we independently align each image’s local pointmap to the global pointmap using ICP and use aligned local pointmaps for evaluation.
@@ 5.1 Scaling the number of views
Fast3R is able to use all-to-all attention during training, which lets it learn from the global context. We hypothesize that the additional context provided by more views during training allows the model to learn higher-order correspondences between multiple frames, ultimately increasing model performance and increasing potential for scaling.
Figures 6 and 7 show that training on increasingly more views consistently improves RRA and RTA for visual odometry, and reconstruction accuracy—even when the number of views used during evaluation is held constant and the model is ultimately evaluated on fewer views than were seen during training. We further evaluate the model’s ability to reason about additional views by increasing the number of images that Fast3R sees during inference. Figure 4 and Figure 5 indicate that as the model uses more views, the average per-view performance improves. This behavior holds for all evaluated metrics in both camera pose estimation and reconstruction. As shown in Figure 5, the model has a better per-view accuracy using 50 images than it does with 20, even though it was trained with 20. Many applications (e.g. reconstruction, odometry) require inference on many views, which is a major motivation for Fast3R removing the pairwise constraint.
5.2 Model scaling and data scaling
The Transformer architecture’s scalability is a key advantage of Fast3R. Although full experiments are in the Appendix A and B, our results show that increasing model size and data consistently boosts performance, promising better outcomes with more computational investment.
@@ 5.3 Training without position interpolation
In Section 3.3, we introduced a randomized version of [5] to enable inference on more views than seen training. Without this technique, model accuracy quickly degrades for pointmap corresponding to image indexes outside the training range, as shown in Figure 8 (top). A version of Fast3R trained on $N=4$ views still produces high-quality pointmaps for views in slot 5 to 24 (Figure 8 bottom).
Pointmap Type 7-Scenes [47] NRGBD [3] DTU
@@
local aligned to global 2.84 1.37 4.39 1.28 3.91 1.41
global 4.81 1.64 4.85 1.32 3.88 1.41
@@
5.4 Inference without local head
In Table 5, we perform an ablation experiment on the effect of using local vs. global pointmaps for 3D reconstruction (also see the visualization in Figure 14 in appendix). Specifically, we compare doing 3D reconstruction directly with the predicted global pointmaps to reconstruction with the predicted local pointmaps aligned to the global coordinate system using ICP (Sec. 4.3). Qualitative and quantitative results show that the local head produces more accurate pointmaps (fewer floaters, less smearing, less distortion) compared to the global head. We attribute this behavior to the local head being more invariant during training than the global head. That is, the local head is supervised such that the 3D XYZ locations of pixels in an image do not change, no matter which view is selected as the anchor view $I_{1}$; whereas for the global head, the XYZ locations for pixels in a view is dependent on which view is selected as the anchor view. Conceptually, the global head needs to learn both 2D-to-3D geometry and the rigid transformation of 3D points in the global coordinate system; whereas the local head only needs to learn 2D-to-3D geometry.
@@ 6 Conclusion
The architecture of Fast3R allows for parallel processing of many views, and its positional embedding design enables “train short, test long” in terms of context length of views. However, we observed that for scenes with very large reconstruction areas, when the number of views becomes extreme (e.g., more than 300), the point map of some views (particularly those with a low confidence score) begins to exhibit drifting behavior. One current way to address this issue is to drop frames with low confidence scores. In dense reconstruction, this approach typically does not hurt reconstruction quality too much. However, to fundamentally address this problem, we hypothesize that future work could explore the following avenues: (1) incorporating more data containing large scenes to improve generalization to such cases; (2) designing better positional embeddings inspired by state-of-the-art long-context language models [76], which can handle very long context lengths and exploit the temporal structure of ordered image sequences (e.g., video).